Hogan Industries had the following transactions occur during 2014,

February 1, 2014 - 36 units at $45 per unit

March 14, 2014 - 62 units at $47 per unit

May 1, 2014 - 44 units at $49 per unit

The company sold 102 units at $63 each and has tax rate of 30%. Assuming that a periodic inventory sustem is used, what is the company's gross profit using FIFO?
Answer #1
Calculation of Gross Profit
Cost Inventory by date
Date Units Per Unit Cost Total cost allocation Inventory Value for allocated Inventory Descreption
Feb' 1 2014 36 $                  45 $        1,620                                                              36                                                   1,620 all units are sold Considering FIFO method
March' 14 2014 62 $                  47 $        2,914                                                              62                                                   2,914 all units are sold Considering FIFO method
May 1 2014 44 $                  49 $        2,156                                                                4                                                       196 Total Units sold is 103 Less feb 1units of 36 & march 14 62 units are used ((103-36-62=4)*49$ = 196$)
TOTAL Cost of Units sold(A) $                                         4,730.00
Totals Sales Value (102*63$) (B) $                                         6,426.00
Gross Profit (B-A) $                                         1,696.00